
None.
| Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
|---|---|---|---|---|---|---|
| 2021-10-07 | 0.0 | 2.0 | 0.0 | 8.25 | 0.0 | 10.25 |
Links
[1] mailto:alistair.walker@noirlab.edu
[2] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/39850/delete?destination=print/node/18206
[3] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/39850/edit?destination=print/node/18206
[4] mailto:alfredo.zenteno@noirlab.edu
[5] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/39851/delete?destination=print/node/18206
[6] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/39851/edit?destination=print/node/18206
[7] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/39854/delete?destination=print/node/18206
[8] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/39854/edit?destination=print/node/18206
[9] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/add/node/18206?destination=print/node/18206
First half: Engineering. DECal scan for filter N419 completed with no problems. Dome opened at 9:48pm, just to be closed after 15 minutes due to rapid increase of humidity.
Second half: Didnt observe. dome closed.